When people smoke, the nicotine they absorb converts to cotinine, which can be measured. A sample of 40 smokers has a mean cotinine level of 172.5. Assuming that the population deviation of cotinine levels is known to be 119.5, find an estimate of the 90% confidence interval from the mean cotinine level for all smokers.

a) 160.25 < µ < 190.28

b) 141.42 < µ < 198.23

c) 138.24 < µ < 210.25

d) 141.42 < µ < 203.58
Answer #1

Population mean, \mu = 172.5

Standard deviation, \sigma = 119.5

Sample size, n = 40

According to Central Limit Theorem, the distribution of sample mean, T is approximately normal.

Confidence interval, CI = \mu \pm Z* \sigma/\sqrt{n}

\sigma_\bar{x} = \sigma/\sqrt{n}

= 119.5/\sqrt{40}

= 18.895

For 90% confidence level, Z* = 1.645

CI = 172.5 \pm 1.645 x 18.895

= 172.5 \pm 31.08

= 141.42 < \mu < 203.58